# Stability via resampling: statistical problems beyond the real line

### v1

@@ 6.2 Results
$\delta=\frac{1}{n}\sum_{i=1}^{n}1_{\|\hat{w}-\hat{w}^{\setminus i}\|>\varepsilon}.$
@@
Spectral analysis (LSSA) $40.033$ $0.3017$
Softmax $0.0744$ $0.0014$
@@ Appendix A Proofs for the Hilbert space setting
$\leq\frac{\mathsf{rad}_{\mathcal{H}}(\mathcal{W})}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2},$
with the last step holding because we have
@@
where the first equality holds since $\hat{w}=\mathbb{E}[\hat{w}^{r}]$.
@@ B.1 Proof of main theorem
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot A+\frac{A^{2}}{3}\ \Longrightarrow\ \sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq A^{2}\cdot\left(\frac{1+\sqrt{7/3}}{2}\right)^{2}.$
@@ B.2 Is 2 needed?
Let $\mathcal{Z}=\mathbb{N}$, the set of positive integers. Let $\mathcal{B}=L_{1}(\mathbb{N})$ (equipped with the norm $\|x\|=\frac{1}{2}\sum_{i\geq 1}|x_{i}|$), and let $\mathcal{W}=\Delta_{\infty}$. Then there exists a (deterministic) algorithm $\mathcal{A}:\cup_{n\geq 0}\mathcal{Z}^{n}\times[0,1]\to\mathcal{W}$, such that for any distribution $\mathcal{Q}_{n}$ satisfying 1, derandomized bagging fails to satisfy mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ for any $\beta<p$.
@@
the set of values appearing in the data set. Then define the algorithm as $\mathcal{A}(\mathcal{D};\xi)=\mathbf{e}_{\psi(S(\mathcal{D}))}\in\mathcal{W},$ where $\mathbf{e}_{j}$ is the $j$th canonical basis vector, with a $1$ in the $j$th location and 00’s elsewhere.
@@
since $\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in(S(\mathcal{D}^{r}))\}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in r\}=p_{\mathcal{Q}_{n}}$. Since this holds for every $i$, we have $\frac{1}{n}\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\backslash i}\|^{2}=p_{\mathcal{Q}_{n}}^{2}$, which completes the proof. ∎
@@ B.3 Additional results
$\delta\epsilon^{2}\geq 1.6\left[R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}+2\rho\cdot p_{\mathcal{Q}_{n}}\right]^{2}.$
@@
$\beta^{2}:=\inf_{\varepsilon>0}\Bigg{\{}(1+\epsilon)\cdot 1.6\bigg{[}R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\\ {}+2\rho\cdot p_{\mathcal{Q}_{n}}\bigg{]}^{2}{}+(1+\epsilon^{-1})\cdot\left(\frac{8KR^{2}}{B}+8\rho^{2}\right)\Bigg{\}}$
@@
$\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\right]=\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})+(\tilde{w}-\tilde{w}^{\setminus i})\|^{2}\right]\\ \leq\frac{2(1+\epsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}\right]+\frac{2(1+\epsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}\right]+\frac{1+\epsilon}{n}\sum_{i\in[n]}\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2},$
since $(a+b+c)^{2}\leq(1+\epsilon)a^{2}+(1+\epsilon^{-1})(b+c)^{2}\leq(1+\epsilon)a^{2}+2(1+\epsilon^{-1})(b^{2}+c^{2})$, for any $a,b,c\geq 0$. The last term in the upper bound is bounded by Theorem 10. Now we bound the remaining terms.

### v2

@@ 6.2 Results
$\delta=\frac{1}{n}\sum_{i=1}^{n}1_{\|\hat{w}-\hat{w}^{\setminus i}\|\geq\varepsilon}.$
@@
Spectral analysis (LSSA) $40.033$ $0.0793$
Softmax $0.0186$ $0.0006$
@@ Appendix A Proofs for the Hilbert space setting
$\leq\frac{\mathsf{rad}_{\mathcal{H}}(\mathcal{W})}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2}.$
The last step holds because we have
@@
where the first equality holds since $\hat{w}=\mathbb{E}[\hat{w}^{r}]$ minimizes mean-square error.
@@ B.1 Proof of main theorem
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot A+\frac{A^{2}}{3}\ \Longrightarrow\ \sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq A^{2}\cdot\left(\frac{1+\sqrt{7/3}}{2}\right)^{2}\leq 1.6A^{2}.$
@@ B.2 Is 2 needed?
Let $\mathcal{Z}=\mathbb{N}$, the set of positive integers. Let $\mathcal{B}=L_{1}(\mathbb{N})$ (equipped with the norm $\|x\|=\frac{1}{2}\sum_{i\geq 1}|x_{i}|$), and let $\mathcal{W}=\Delta_{\infty}$. Then there exists a (deterministic) algorithm $\mathcal{A}:\cup_{n\geq 0}\mathcal{Z}^{n}\times[0,1]\to\mathcal{W}$, such that for all $n\geq 1$, for any distribution $\mathcal{Q}_{n}$ satisfying 1, derandomized bagging fails to satisfy mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ for any $\beta<p_{\mathcal{Q}_{n}}$.
@@
the set of unique values appearing in the data set. Then define $\mathcal{A}(\mathcal{D};\xi)=\mathbf{e}_{\psi(S(\mathcal{D}))}\in\mathcal{W},$ where $\mathbf{e}_{j}$ is the $j$th canonical basis vector, with a $1$ in the $j$th location and 00’s elsewhere.
@@
since $\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in S(\mathcal{D}^{r})\}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in r\}=p_{\mathcal{Q}_{n}}$. Since this holds for every $i$, we have $\frac{1}{n}\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\backslash i}\|^{2}=p_{\mathcal{Q}_{n}}^{2}$, which completes the proof. ∎
@@ B.3 Additional results
$\delta\varepsilon^{2}\geq 1.6\left[R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}+2\rho\cdot p_{\mathcal{Q}_{n}}\right]^{2}.$
@@
[…]
$\beta^{2}:=\inf_{\varepsilon>0}\Bigg{\{}(1+\varepsilon)\cdot 1.6\bigg{[}R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\\ {}+2\rho\cdot p_{\mathcal{Q}_{n}}\bigg{]}^{2}{}+(1+\varepsilon^{-1})\cdot\left(\frac{8KR^{2}}{B}+8\rho^{2}\right)\Bigg{\}}$
@@
$\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\right]=\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})+(\tilde{w}-\tilde{w}^{\setminus i})\|^{2}\right]\\ \leq\frac{1+\varepsilon}{n}\sum_{i\in[n]}\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2}+\frac{2(1+\varepsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}\right]+\frac{2(1+\varepsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}\right],$
[…]